MULTIPLE PATHWAYS IN THE VIBRATIONAL AUTOIONIZATION OF THE BENDING OVERTONES OF FORMYL RADICAL.

\underline{HARTMUT G. HEDDERICH}, ERIC J. ZÜCKERMAN and EDWARD R. GRANT, Department of Chemistry, Purdue University, West Lafayette, IN 47907-1393.

Selected individual rotational levels in vibrational bands of the Renner-Teller split (020) and (030) systems in the 3pπ^{2}Π Rydberg state of HCO have served as originating levels for vertical second-photon transitions to vibrationally autoionizing Rydberg series. Spectra of individual rotational levels associated with the (02^{0}0) and (02^{2}0) as well as the (03^{1}0) and (03^{3}0) states of HCO^{+} were obtained. Linewidths of assigned series convey information on autoionization dynamics. Series extrapolations determine anharmonic terms in the HCO^{+} bending potential.